# Optical aberrations in autonomous driving: Physics-informed parameterized temperature scaling for neural network uncertainty calibration

### v1

@@ Abstract
A trustworthy representation of uncertainty is desirable and should be considered as a key feature of any machine learning method (Huellermeier and Waegeman,, 2021). This conclusion of Huellermeier et al. underpins the importance of calibrated uncertainties. Since AI-based algorithms are heavily impacted by dataset shifts, the automotive industry needs to safeguard its system against all possible contingencies. One important but often neglected dataset shift is caused by optical aberrations induced by the windshield. For the verification of the perception system performance, requirements on the AI performance need to be translated into optical metrics by a bijective mapping (Braun,, 2023). Given this bijective mapping it is evident that the optical system characteristics add additional information about the magnitude of the dataset shift. As a consequence, we propose to incorporate a physical inductive bias into the neural network calibration architecture to enhance the robustness and the trustworthiness of the AI target application, which we demonstrate by using a semantic segmentation task as an example. By utilizing the Zernike coefficient vector of the optical system as a physical prior we can significantly reduce the mean expected calibration error in case of optical aberrations. As a result, we pave the way for a trustworthy uncertainty representation and for a holistic verification strategy of the perception chain.
Keywords:
@@
Neural Network Calibration Predictive Uncertainty Dataset shifts Wavefront aberrations Autonomous driving.1 Introduction
@@
Autonomously driving cars perceive the environment through different sensor, e.g., wide-view cameras, telephoto cameras, Light Detection and Ranging (LiDAR) sensors etc. Typically, the measured sensor signals serve as the input to a neural network, which is supposed to predict the actions required (e.g., acceleration, steering angle etc.) to reach the next state. In the development phase, the neural network is trained on a training dataset that is typically captured by a small fleet of test mules. If everything goes well with the architectural design and the neural network demonstrates sufficient accuracy and generalizability, then the autonomous driving functionality might be considered as operational. As the conceptual phase is taken to serial production, the real-world performance of the AI-based driving function might differ by a huge margin from what has been observed during development. A prominent contributor to this phenomenon is the perception chain of the telephoto camera (i.e., a camera with a telephoto lens). Telephoto cameras are distinguished by a long focal length, which results in a high pixel resolution per field-angle. As a consequence, the target application of telephoto cameras is object detection and classification for far-field objects, especially important on highways with high driving speeds. Unfortunately, this benefit also comes with a downside, namely an increased sensitivity for optical aberrations. Every car has a windshield, and every windshield has its unique aberration pattern. This has not been an issue for standard automotive cameras because the width of the blurring kernel induced by the windshield was always smaller than a pixel-pitch of the Complementary Metal-Oxide-Semiconductor (CMOS) sensor. With the use of telephoto cameras, this does not hold true anymore and the images captured might by heavily impacted by the windshield in terms of sharpness. This gives rise to a shift in image quality between the test mule recordings used for training and the car-by-car perception chain.
@@
Dataset shifts are of major concern for the homologation of safety-critical autonomous driving functions. Dataset shifts are generally given if the network infers information from an instance, which does not share the same underlying probability density function (PDF) as the training dataset distribution. Consequently, the neural network utilizes the learned functional relationship between input and output for extrapolating into a different domain. This gives rise to a performance drop of the model. The performance is not only affected in terms of the target key performance indicator (KPI) but also the corresponding uncertainty estimation might become biased (Wolf et al., 2023c, ).
@@
According to information theory, the total predictive uncertainty for a classification problem is given by the Shannon entropy (Huellermeier and Waegeman,, 2021). From a metrological perspective, if the model is properly calibrated, the uncertainty measure is expected to align with the observed error rate in the network’s predictions. Equivalently speaking, a perfectly calibrated network is given if the confidence estimate is congruent to the measured prediction accuracy. For assessing the calibration performance, there exists an entire zoo of measures, e.g., the Uncertainty Calibration Error (UCE) (Pakdaman Naeini et al.,, 2015), the Area Under the Sparsification Error curve (AUSES) (Dreissig et al.,, 2023) utilizing the Shannon entropy (Shannon,, 1948) for sorting, the Expected Calibration Error (ECE) (Pakdaman Naeini et al.,, 2015) or the AUSEV utilizing the variation ratio (Maag et al.,, 2020) for sorting. The conceptual differences between point-wise predictive uncertainty calibration estimators (ECE, AUSEV) and entropy-based calibration measures (UCE, AUSES) results in a decoupling of neural network calibration measures (Wolf et al., 2024a, ). As a consequence, it is essential to make a physical sound decision on which calibration measure to employ for the neural network under consideration.
@@
In this paper, we focus on semantic segmentation because of the hypothesis that a dataset shift in terms of sharpness will affect a pixel-wise prediction the most. At this point, it is important to underscore that, irrespective of the specific AI task selected, this methodology exhibits encouraging potential for effective generalization across a broad spectrum of tasks. Our semantic segmentation Convolutional Neural Network (CNN) will employ a negative log-likelihood loss, which makes it favorable to rely on point-wise predictive uncertainty calibration estimators because it matches the nature of the ground truth label distribution, which allocates all statistical mass to the ground truth class and analytically resembles the Kronecker delta function. As a consequence, the expected Shannon information (Shannon,, 1948) is given by the negative logarithm of the probability score predicted by the CNN for the ground truth class and the cross-entropy is minimized by maximizing the prediction confidence for the ground truth class during training. As a result, the expected negative log-likelihood loss, aka. cross-entropy, is invariant under differences in the probability mass allocation over the remaining wrong classes. The same does not hold true for the Shannon entropy (Shannon,, 1948), which is the standard measure for the total predictive uncertainty according to information theory (Huellermeier and Waegeman,, 2021). This gives rise to a degree of freedom in the uncertainty evaluation or to put it differently, the entropy-based uncertainties for two independent instances might differ even though the model confidence predictions are equivalent (Wolf et al., 2024a, ). We avoid this by employing the variation ratio (Maag et al.,, 2020) as a point-wise measure for the prediction uncertainty and the ECE as a point-wise calibration measure.
@@
Calibrated uncertainties are an essential requirement for a physically sound sensor fusion process and for system monitoring. On the one hand, fusing feature attributes should incorporate the associated embedding uncertainties in order to achieve the most reliable latent space representation. On the other hand, in order to safeguard autonomous systems, the prediction uncertainties need to be tracked. If the uncertainty is low, and hence the situation is identified with sufficient confidence, a reliable decision can be made. If the confidence is insufficiently low then an independent secondary system must contribute additional information for the decision-making process or the system has to fall back into a safe state mode automatically. Consequently, the trustworthiness of the uncertainty estimates, quantified by the ECE, is decisive for the reliability of autonomous driving systems. However, if a dataset shift is induced, e.g. by optical aberrations of the windshield, the calibration of the network confidences - and uncertainties vice versa - breaks down and the network becomes increasingly overconfident (Wolf et al., 2023c, ).
@@
To tackle this task, we present a novel neural network architecture, which extends the state-of-the-art Parameterized Temperature Scaling (PTS) (Tomani et al.,, 2022) approach by incorporating a physical inductive bias (Banerjee et al.,, 2024). We demonstrate the benefits of integrating physical priors to PTS, which we will refer to as Physics Informed Parameterized Temperature Scaling (PIPTS), by comparing the results to the state-of-the-art PTS method and to the standard Temperature Scaling (TS) (Guo et al.,, 2017) technique. The physical prior consists of the predicted Zernike coefficient vector of the optical system and is intended to enhance the resilience of the autonomous driving perception pipeline against optical perturbations.
@@
2 Related work
Fundamentally, the temperature determines the sensitivity of the entropy w.r.t. changes in the internal energy from a physical perspective (Goodstein,, 1975) and w.r.t. changes in the logits from an AI point of view (Shannon,, 1948).
@@ 2 Related work
In AI language models, temperature controls the variability in word predictions (Xie et al.,, 2024). The model predicts a likelihood for each word within the vocabulary and the subsequent word is chosen randomly according to the predicted probability mass function. Calibrating the model with a temperature of zero leads to deterministic behavior and complete repeatability. Increasing the temperature allows words with lower likelihoods to still have a chance, creating more diverse results. This process mirrors the Boltzmann distribution in physics, where states are sampled based on energy levels. In AI, the Softmax function performs a similar task, treating model logits as negative energies. Higher temperatures broaden the probability distribution, making all outcomes more equal, while lower temperatures sharpen the probability distribution.
@@
Finding the optimal temperature is the key for establishing model trustworthiness. The straightforward way to determine the temperature is given by minimizing the negative log-likelihood (Guo et al.,, 2017). Unfortunately, this standard Temperature Scaling (TS) methodology is highly limited in terms of the model information capacity (one degree of freedom).
@@
As an extension, Ensemble Temperature Scaling (ETS) (Zhang et al.,, 2020) computes an weighted average over three different calibration maps, the TS calibrator with adjustable temperature , TS with (identity mapping) and TS with (uniform mapping). Hence, ETS has four degrees of freedom.
@@
In order to further increase the information capacity of the calibration method, Parameterized Temperature Scaling (PTS) (Tomani et al.,, 2022) was proposed. PTS leverages a neural network to predict an instance-wise temperature based on the corresponding logit tensor, while preserving model accuracy.
@@
A similar methodology is Sample-Dependent Adaptive Temperature Scaling (Joy et al.,, 2023), which also predicts an instance-wise temperature. In contrast to PTS, the approach leverages the latent space representation of a Variational Autoencoder (VAE) (Kingma and Welling,, 2013) as the input for a post-hoc Multi-Layer Perceptron (MLP) for predicting an instance-wise temperature. The benefit of using the VAE’s latent space embeddings instead of the logit tensor for the post-hoc MLP lies in the effectiveness of the VAE to cluster the predictions based on their calibration quality, which improves the calibration performance of the MLP under distribution shifts (Joy et al.,, 2023).
@@
Most recently, Adaptive Temperature Scaling (ATS) (Krumpl et al.,, 2024) has been proposed as a calibration technique, which enhances the reliability against out-of-distribution samples without the need of training a post-hoc MLP calibrator. The core idea of ATS lies in computing an instance-wise temperature based on the intermediate layer activations of the baseline neural network. After training, the Cumulative Distribution Function (CDF) of the mean activation for each layer is computed across the entire training dataset. During inference, the layer-wise mean activations are compared to the precomputed CDFs to calculate layer-wise p-values (Rudolph et al.,, 2023; Upton and Cook,, 2008), which are mapped to an instance-wise temperature. In addition to an enhanced calibration performance, low temperatures indicate in-distribution samples, while high temperatures suggest out-of-distribution inputs (Krumpl et al.,, 2024).
@@
For all accuracy preserving calibration methods mentioned so far, it is required to assume an uncertainty estimator. In mathematical terms, calibration quality metrics measure the bias of a predictive uncertainty estimator. Consequently, neural network calibration and predictive uncertainty estimation are two distinct concepts. A perfect predictive uncertainty estimator is unbiased such that the calibration temperature is equal to one. There are several different methods to estimate the predictive uncertainty for semantic segmentation (Gawlikowski et al.,, 2023). A very powerful approach are Deep Ensembles (Lakshminarayanan et al.,, 2016) that use several neural networks of the same architecture in parallel but with different initializations to retrieve various subsamples from the posterior distribution. The mean of this subsample is outputted as the model’s prediction and the standard deviation of the mean serves as the predictive uncertainty estimator. A major downside of this approach is the computational overhead generated during training and inference by the utilization of several models that sample the hypothesis space (Huellermeier and Waegeman,, 2021). This problem has been addressed recently in the work of Landgraf et al., 2024b on Deep Uncertainty Distillation using Ensembles for Semantic Segmentation (DUDES). They propose a student-teacher distillation framework where the Deep Ensemble model, referred to as the teacher, is used to guide a less complex model, known as the student, in estimating the ensemble-based predictive uncertainties. This significantly reduces the inference time because only one forward pass is required and the information about the posterior distribution is distilled into the student neural network, wherefore the calibration quality of the predictive uncertainty estimate is maintained. This concept has been further extended by the student-teacher distillation framework for efficient multi-task uncertainties, referred to as EMUFormer (Landgraf et al., 2024a, ). They employ a Deep Ensemble of transformer-based multi-task networks for semantic segmentation and monocular depth estimation (termed as SegDepthFormer) to evaluate the predictive uncertainty. The backbone of this methodology is the idea of enhancing the generalization capabilities of a neural network by multi-task learning. With the SegDepthFormer architecture they demonstrated that this idea can be transformed to the network calibration and the results indicate less biased predictive uncertainty estimates in terms of the mECE if multi-task learning is employed. Nevertheless, we will select the variation ratio (Maag et al.,, 2020) as a measure for the predictive uncertainty in this work because of the non-existent computational overhead and it’s simplicity, which mainly explains why the variation ratio is widely adapted.
@@
In our use case we know what drives the distribution shift, namely optical aberrations within the perception chain, e.g., the windshield or diverse weather phenomena. As a consequence, it seems natural to incorporate this prior knowledge into the calibration process. In order to do so, the optical aberrations need to be estimated online alongside the target application. The most fundamental way to characterize optical aberrations is given by quantifying the optical path difference map in terms of the Zernike coefficients (Zernike and Stratton,, 1934; Zernike,, 1934; Bhatia and Wolf,, 1954). Utilizing a neural network for predicting the Zernike coefficient vector is a well established approach in astronomy (McGuire et al.,, 1999; Andersen et al.,, 2020) as a way to replace the need for on the fly Shack-Hartmann measurements. The Very Large Telescope (VLT) of the European Southern Observatory (ESO) uses the information about the Zernike coefficients to perform an online correction of the wavefront aberrations induced by the atmosphere (Merkle and Hubin,, 1992). This is realized by adjusting deformable mirrors, which is a technique from adaptive optics (Hampson et al.,, 2021).
@@
The work of Jaiswal et al., (2023) on physics-driven turbulence image restoration with stochastic refinement utilizes the Zernike coefficients to parameterize a physics-based turbulence simulator. By coupling the vision transformer-based (Dosovitskiy et al.,, 2021) image restoration network with the Fourier-optical aberration model during training, they are able to effectively disentangle the stochastic degradation caused by atmospheric turbulence from the underlying image. This enhances the generalizability of the image restoration network across real-world datasets with varying turbulence strength (Jaiswal et al.,, 2023). As a consequence, by incorporating a physical inductive bias to the transformer architecture they effectively reduce the sensitivity of their target application on dataset shifts induced by turbulence-driven optical aberrations.

### v2

@@ Abstract
“A trustworthy representation of uncertainty is desirable and should be considered as a key feature of any machine learning method” [1]. This conclusion of Huellermeier et al. underpins the importance of calibrated uncertainties. Since AI-based algorithms are heavily impacted by dataset shifts, the automotive industry needs to safeguard its system against all possible contingencies. One important but often neglected dataset shift is caused by optical aberrations induced by the windshield. For the verification of the perception system performance, requirements on the AI performance need to be translated into optical metrics by a bijective mapping. Given this bijective mapping it is evident that the optical system characteristics add additional information about the magnitude of the dataset shift. As a consequence, we propose to incorporate a physical inductive bias into the neural network calibration architecture to enhance the robustness and the trustworthiness of the AI target application, which we demonstrate by using a semantic segmentation task as an example. By utilizing the Zernike coefficient vector of the optical system as a physical prior we can significantly reduce the mean expected calibration error in case of optical aberrations. As a result, we pave the way for a trustworthy uncertainty representation and for a holistic verification strategy of the perception chain.
Index Terms:
@@
Neural Network Calibration, Predictive Uncertainty, Dataset shifts, Wavefront aberrations, Autonomous driving.I Introduction
@@
Autonomously driving cars perceive the environment through different sensors, e.g., wide-angle cameras, telephoto cameras [2, 3, 4], Light Detection and Ranging (LiDAR) sensors etc. Typically, the measured sensor signals serve as the input to a neural network, which is supposed to predict the actions required (e.g., acceleration, steering angle etc.) to reach the next state. In the development phase, the neural network is trained on a training dataset that is typically captured by a small fleet of test mules. If everything goes well with the architectural design and the neural network demonstrates sufficient accuracy and generalizability, then the autonomous driving functionality might be considered as operational. As the conceptual phase is taken to serial production, the real-world performance of the AI-based driving function might differ by a huge margin from what has been observed during development. A prominent contributor to this phenomenon is the perception chain of the telephoto camera (i.e., a camera with a telephoto lens). Telephoto cameras are distinguished by a long focal length, which results in a high pixel resolution per field-angle. As a consequence, the target application of telephoto cameras is object detection and classification for far-field objects, especially important on highways with high driving speeds. Unfortunately, this benefit also comes with a downside, namely an increased sensitivity for optical aberrations. Every car has a windshield, and every windshield has its unique aberration pattern. This has not been an issue for standard automotive cameras because the width of the blurring kernel induced by the windshield was always smaller than a pixel-pitch of the Complementary Metal-Oxide-Semiconductor (CMOS) sensor. With the use of telephoto cameras, this does not hold true anymore and the images captured might by heavily impacted by the windshield in terms of sharpness. This gives rise to a shift in image quality between the test mule recordings used for training and the car-by-car perception chain.
@@
Dataset shifts are of major concern for the homologation of safety-critical autonomous driving functions. Dataset shifts are generally given if the network infers information from an instance, which does not share the same underlying probability density function (PDF) as the training dataset distribution. Consequently, the neural network utilizes the learned functional relationship between input and output for extrapolating into a different domain. This gives rise to a performance drop of the model. The performance is not only affected in terms of the target key performance indicator (KPI) but also the corresponding uncertainty estimation might become biased [5].
@@
According to information theory, the total predictive uncertainty for a classification problem is given by the Shannon entropy [1]. From a metrological perspective, if the model is properly calibrated, the uncertainty measure is expected to align with the observed error rate in the network’s predictions. Equivalently speaking, a perfectly calibrated network is given if the confidence estimate is congruent to the measured prediction accuracy. For assessing the calibration performance, there exists an entire zoo of measures, e.g., the Uncertainty Calibration Error (UCE) [6], the Area Under the Sparsification Error curve (AUSES) [7] utilizing the Shannon entropy [8] for sorting, the Expected Calibration Error (ECE) [6] or the AUSEV utilizing the variation ratio [9] for sorting. The conceptual differences between point-wise predictive uncertainty calibration estimators (ECE, AUSEV) and entropy-based calibration measures (UCE, AUSES) results in a decoupling of neural network calibration measures [10]. As a consequence, it is essential to make a physical sound decision on which calibration measure to employ for the neural network under consideration.
@@
In this paper, we focus on semantic segmentation because of the hypothesis that a dataset shift in terms of sharpness will affect a pixel-wise prediction the most. At this point, it is important to underscore that, irrespective of the specific AI task selected, this methodology exhibits encouraging potential for effective generalization across a broad spectrum of tasks. Our semantic segmentation Convolutional Neural Network (CNN) will employ a negative log-likelihood loss, which makes it favorable to rely on point-wise predictive uncertainty calibration estimators because it matches the nature of the ground truth label distribution, which allocates all statistical mass to the ground truth class and analytically resembles the Kronecker delta function. As a consequence, the expected Shannon information [8] is given by the negative logarithm of the probability score predicted by the CNN for the ground truth class and the cross-entropy is minimized by maximizing the prediction confidence for the ground truth class during training. As a result, the expected negative log-likelihood loss, aka. cross-entropy, is invariant under differences in the probability mass allocation over the remaining wrong classes. The same does not hold true for the Shannon entropy [8], which is the standard measure for the total predictive uncertainty according to information theory [1]. This gives rise to a degree of freedom in the uncertainty evaluation or to put it differently, the entropy-based uncertainties for two independent instances might differ even though the model confidence predictions are equivalent [10]. We avoid this by employing the variation ratio [9] as a point-wise measure for the prediction uncertainty and the ECE as a point-wise calibration measure.
@@
Calibrated uncertainties are an essential requirement for a physically sound sensor fusion process and for system monitoring. On the one hand, fusing feature attributes should incorporate the associated embedding uncertainties in order to achieve the most reliable latent space representation. On the other hand, in order to safeguard autonomous systems, the prediction uncertainties need to be tracked. If the uncertainty is low, and hence the situation is identified with sufficient confidence, a reliable decision can be made. If the confidence is insufficiently low then an independent secondary system must contribute additional information for the decision-making process or the system has to fall back into a safe state mode automatically. Consequently, the trustworthiness of the uncertainty estimates, quantified by the ECE, is decisive for the reliability of autonomous driving systems. However, if a dataset shift is induced, e.g. by optical aberrations of the windshield, the calibration of the network confidences - and uncertainties vice versa - breaks down and the network becomes increasingly overconfident [5].
@@
To tackle this task, we present a novel neural network architecture, which extends the state-of-the-art Parameterized Temperature Scaling (PTS) [11] approach by incorporating a physical inductive bias [12]. We demonstrate the benefits of integrating physical priors to PTS, which we will refer to as Physics Informed Parameterized Temperature Scaling (PIPTS), by comparing the results to the state-of-the-art PTS method and to the standard Temperature Scaling (TS) [13] technique. The physical prior consists of the predicted Zernike coefficient vector of the optical system and is intended to enhance the resilience of the autonomous driving perception pipeline against optical perturbations.
@@
II Related work
Fundamentally, the temperature determines the sensitivity of the entropy w.r.t. changes in the internal energy from a physical perspective [14] and w.r.t. changes in the logits from an AI point of view [8].
@@ II Related work
In AI language models, temperature controls the variability in word predictions [15]. The model predicts a likelihood for each word within the vocabulary and the subsequent word is chosen randomly according to the predicted probability mass function. Calibrating the model with a temperature of zero leads to deterministic behavior and complete repeatability. Increasing the temperature allows words with lower likelihoods to still have a chance, creating more diverse results. This process mirrors the Boltzmann distribution in physics, where states are sampled based on energy levels. In AI, the Softmax function performs a similar task, treating model logits as negative energies. Higher temperatures broaden the probability distribution, making all outcomes more equal, while lower temperatures sharpen the probability distribution.
@@
Finding the optimal temperature is the key for establishing model trustworthiness. The straightforward way to determine the temperature is given by minimizing the negative log-likelihood [13]. Unfortunately, this standard Temperature Scaling (TS) methodology is highly limited in terms of the model information capacity (one degree of freedom).
@@
As an extension, Ensemble Temperature Scaling (ETS) [16] computes an weighted average over three different calibration maps, the TS calibrator with adjustable temperature , TS with (identity mapping) and TS with (uniform mapping). Hence, ETS has four degrees of freedom.
@@
In order to further increase the information capacity of the calibration method, Parameterized Temperature Scaling (PTS) [11] was proposed. PTS leverages a neural network to predict an instance-wise temperature based on the corresponding logit tensor, while preserving model accuracy.
@@
A similar methodology is Sample-Dependent Adaptive Temperature Scaling [17], which also predicts an instance-wise temperature. In contrast to PTS, the approach leverages the latent space representation of a Variational Autoencoder (VAE) [18] as the input for a post-hoc Multi-Layer Perceptron (MLP) for predicting an instance-wise temperature. The benefit of using the VAE’s latent space embeddings instead of the logit tensor for the post-hoc MLP lies in the effectiveness of the VAE to cluster the predictions based on their calibration quality, which improves the calibration performance of the MLP under distribution shifts [17].
@@
Most recently, Adaptive Temperature Scaling (ATS) [19] has been proposed as a calibration technique, which enhances the reliability against out-of-distribution samples without the need of training a post-hoc MLP calibrator. The core idea of ATS lies in computing an instance-wise temperature based on the intermediate layer activations of the baseline neural network. After training, the Cumulative Distribution Function (CDF) of the mean activation for each layer is computed across the entire training dataset. During inference, the layer-wise mean activations are compared to the precomputed CDFs to calculate layer-wise p-values [20, 21], which are mapped to an instance-wise temperature. In addition to an enhanced calibration performance, low temperatures indicate in-distribution samples, while high temperatures suggest out-of-distribution inputs [19].
@@
For all accuracy preserving calibration methods mentioned so far, it is required to assume an uncertainty estimator. In mathematical terms, calibration quality metrics measure the bias of a predictive uncertainty estimator. Consequently, neural network calibration and predictive uncertainty estimation are two distinct concepts. A perfect predictive uncertainty estimator is unbiased such that the calibration temperature is equal to one. There are several different methods to estimate the predictive uncertainty for semantic segmentation [22]. A very powerful approach are Deep Ensembles [23] that use several neural networks of the same architecture in parallel but with different initializations to retrieve various subsamples from the posterior distribution. The mean of this subsample is outputted as the model’s prediction and the standard deviation of the mean serves as the predictive uncertainty estimator. A major downside of this approach is the computational overhead generated during training and inference by the utilization of several models that sample the hypothesis space [1]. This problem has been addressed recently in the work of Landgraf et al. on Deep Uncertainty Distillation using Ensembles for Semantic Segmentation (DUDES) [24]. They propose a student-teacher distillation framework where the Deep Ensemble model, referred to as the teacher, is used to guide a less complex model, known as the student, in estimating the ensemble-based predictive uncertainties. This significantly reduces the inference time because only one forward pass is required and the information about the posterior distribution is distilled into the student neural network, wherefore the calibration quality of the predictive uncertainty estimate is maintained. This concept has been further extended by the student-teacher distillation framework for efficient multi-task uncertainties, referred to as EMUFormer [25]. They employ a Deep Ensemble of transformer-based multi-task networks for semantic segmentation and monocular depth estimation (termed as SegDepthFormer) to evaluate the predictive uncertainty. The backbone of this methodology is the idea of enhancing the generalization capabilities of a neural network by multi-task learning. With the SegDepthFormer architecture they demonstrated that this idea can be transformed to the network calibration and the results indicate less biased predictive uncertainty estimates in terms of the mECE if multi-task learning is employed. Nevertheless, we will select the variation ratio [9] as a measure for the predictive uncertainty in this work because of the non-existent computational overhead and it’s simplicity, which mainly explains why the variation ratio is widely adapted.
@@
In our use case we know what drives the distribution shift, namely optical aberrations within the perception chain, e.g., the windshield or diverse weather phenomena. As a consequence, it seems natural to incorporate this prior knowledge into the calibration process. In order to do so, the optical aberrations need to be estimated online alongside the target application. The most fundamental way to characterize optical aberrations is given by quantifying the optical path difference map in terms of the Zernike coefficients [26, 27, 28]. Utilizing a neural network for predicting the Zernike coefficient vector is a well established approach in astronomy [29, 30] as a way to replace the need for on the fly Shack-Hartmann measurements. The Very Large Telescope (VLT) of the European Southern Observatory (ESO) uses the information about the Zernike coefficients to perform an online correction of the wavefront aberrations induced by the atmosphere [31]. This is realized by adjusting deformable mirrors, which is a technique from adaptive optics [32].
@@
The work of Jaiswal et al. on physics-driven turbulence image restoration with stochastic refinement [33] utilizes the Zernike coefficients to parameterize a physics-based turbulence simulator. By coupling the vision transformer-based [34] image restoration network with the Fourier-optical aberration model during training, they are able to effectively disentangle the stochastic degradation caused by atmospheric turbulence from the underlying image. This enhances the generalizability of the image restoration network across real-world datasets with varying turbulence strength [33]. As a consequence, by incorporating a physical inductive bias to the transformer architecture they effectively reduce the sensitivity of their target application on dataset shifts induced by turbulence-driven optical aberrations.